OpenStudy (anonymous):

Let f(x) = x^2 - 16. Find f-1(x).

OpenStudy (lgbasallote):

first step: change f(x) into y \[\implies y = x^2 - 16\] now do you know how to isolate x in this equation?

hero (hero):

Hint set y = f(x) to get y = x^2 - 16 Swap x and y to get x = y^2 - 16 Now isolate and solve for y Finally replace the y variable with f^{-1}(x) after isolation

OpenStudy (anonymous):

i'm stuck at the step where its √x = √y^2 - 16 i know the x becomes x^2 but i don't know what to do with the other side. and from there.

hero (hero):

Well, for one, you forgot to swap x and y. Secondly, you're supposed to add 16 to both sides afterwards. then you take the square root.

hero (hero):

Now all you have to do is remember this general rule \(x = \pm \sqrt{a}\)

OpenStudy (anonymous):

so would it be \[\pm 4 \sqrt{x}\] ?

hero (hero):

no, you didn't follow the steps properly

hero (hero):

You cannot reduce the square root of an addition

OpenStudy (anonymous):

\[\pm \sqrt{x + 16}\] then?

hero (hero):

good job

hero (hero):

Don't forget to set that equal to the inverse of the function